# TAPTRv3: Spatial and Temporal Context Foster Robust Tracking of Any Point in Long Video

### v1

@@ A.2 More Implementation Details
A.3 More Visualizations
@@
More Visual Comparison with TAPTRv2. We provide more visualizations of the comparisons between TAPTRv2 and TAPTRv3, demonstrating the superiority of TAPTRv3. For more details, please refer Fig. 10, Fig. 9, Fig. 11, and their corresponding captions.
@@
Figure 9: Visual comparison between TAPTRv2 and TAPTRv3. Best view in electronic version. From the third image in the first row (36th frame), TAPTRv2 loses tracking of the turtle’s flippers, and in the last few frames loses tracking of the turtle shell and the point on the fish below the turtle. TAPTRv3, on the other hand, maintains stable and accurate tracking throughout the video.
Figure 10: Visual comparison between TAPTRv2 and TAPTRv3. Best view in electronic version. When the goldfish is about to swim out of the frame from right to left (119th frame), TAPTRv2 loses many target tracking points. Afterward, the goldfish swims back from left to right, and starting from the 358th frame, the video shows the other side of the goldfish, where the original target tracking points are occluded. However, TAPTRv2 incorrectly estimates them as visible or on another fish. TAPTRv3, on the other hand, maintains the correct estimation. Until the last dozens of frames, when the goldfish turns around again, TAPTRv3 successfully detects the initial target tracking points, estimates them as visible, and provides accurate positions.
Figure 11: Visual comparison between TAPTRv2 and TAPTRv3. Best view in electronic version. Over time, TAPTRv2 incorrectly estimates the location and visibility of points on jellyfish, and the error accumulates, while TAPTRv3’s results are more accurate.
Figure 12: Visual comparison between TAPTRv3 with and without the auto-triggered global matching. After the occluder appears and then disappears, TAPTRv3 without auto-triggered global matching takes about 70 frames to successfully re-track the target tracking points. However, with the help of global matching, this process takes only two frames.
Figure 13: Visual comparison between TAPTRv3 with and without the auto-triggered global matching. After the occluder appears and then disappears, TAPTRv3 without auto-triggered global matching takes about 14 frames to successfully re-track the target tracking points. However, with the help of global matching, this process takes only two frames.
Figure 14: Additional visualizations of TAPTRv3’s robust predictions.

### v2

@@ A.4 More Inplementation Details
A.5 Limitations
TAPTRv3 is an online tracker that can process streaming video input, meaning that the decoder handles only one frame at a time. This design leads to very low computational cost (see Sec. A.2) but comes with a limitation in parallelism. Even with sufficient GPU memory, the model can not fully utilize it. As a result, TAPTRv3 may appear slower than offline methods when evaluated in offline settings, since those methods can process multiple frames simultaneously in a single forward pass, while TAPTRv3 requires multiple passes. However, TAPTRv3 can still be adapted for offline use by simply applying a sliding window strategy with an increased window size to improve parallelism. While this adjustment may incur a slight performance drop, it provides a flexible trade-off between speed and accuracy depending on the application scenario.
A.6 More Visualizations
@@
More Visual Comparison with TAPTRv2. We provide more visualizations of the comparisons between TAPTRv2 and TAPTRv3, demonstrating the superiority of TAPTRv3. For more details, please refer to Fig. 8, Fig. 9, Fig. 10, and their corresponding captions.
@@
Figure 8: Visual comparison between TAPTRv2 and TAPTRv3. Best view in electronic version. From the third image in the first row (36th frame), TAPTRv2 loses tracking of the turtle’s flippers, and in the last few frames loses tracking of the turtle shell and the point on the fish below the turtle. TAPTRv3, on the other hand, maintains stable and accurate tracking throughout the video. The corresponding videos (CompareVideo1_\_TAPTRv2.mp4 and CompareVideo1_\_TAPTRv3.mp4) are provided in the supplementary material.
Figure 9: Visual comparison between TAPTRv2 and TAPTRv3. When the goldfish is about to swim out of the frame from right to left (119th frame), TAPTRv2 loses many target tracking points. Afterward, the goldfish swims back from left to right, and starting from the 358th frame, the video shows the other side of the goldfish, where the original target tracking points are occluded. However, TAPTRv2 incorrectly estimates them as visible or on another fish. TAPTRv3, on the other hand, maintains the correct estimation. Until the last dozens of frames, when the goldfish turns around again, TAPTRv3 successfully detects the initial target tracking points, estimates them as visible, and provides accurate positions. The corresponding videos (CompareVideo2_\_TAPTRv2.mp4 and CompareVideo2_\_TAPTRv3.mp4) are provided in the supplementary material.
Figure 10: Visual comparison between TAPTRv2 and TAPTRv3. Best view in electronic version. Over time, TAPTRv2 incorrectly estimates the location and visibility of points on jellyfish, and the error accumulates, while TAPTRv3’s results are more accurate. The corresponding videos (CompareVideo3_\_TAPTRv2.mp4 and CompareVideo3_\_TAPTRv3.mp4) are provided in the supplementary material.
Figure 11: Visual comparison between TAPTRv3 with and without the auto-triggered global matching. After the occluder appears and then disappears, TAPTRv3 without auto-triggered global matching takes about 70 frames to successfully re-track the target tracking points. However, with the help of global matching, this process takes only two frames. The corresponding videos (CompareVideo4_\_TAPTRv3wGM.mp4 and CompareVideo4_\_TAPTRv3woGM.mp4) are provided in the supplementary material.
Figure 12: Visual comparison between TAPTRv3 with and without the auto-triggered global matching. After the occluder appears and then disappears, TAPTRv3 without auto-triggered global matching takes about 14 frames to successfully re-track the target tracking points. However, with the help of global matching, this process takes only two frames. The corresponding videos (CompareVideo5_\_TAPTRv3wGM.mp4 and CompareVideo5_\_TAPTRv3woGM.mp4) are provided in the supplementary material.
Figure 13: Additional visualizations of TAPTRv3’s robust predictions. The corresponding videos (RobustVideo1.mp4 and RobustVideo2.mp4) are provided in the supplementary material.